OpenStudy (anonymous):

A bowling ball slows down at a rate of .12 m/s^2. 5 seconds after it starts down the bowling alley its speed is 15m/s. what was its initial speed?

OpenStudy (danjs):

Vf = Vi +a*t

OpenStudy (danjs):

15 = Vi + (-0.12)(5)

OpenStudy (danjs):

Initial Velocity = 15 + 0.12*5 seconds

OpenStudy (danjs):

15.6